Please see inventory related information for Dell and HP in the year 2012 below:

Types of inventory

Dell

HP

Cost of Goods Sold

$10,105

$21,820

Production materials (PARTS)

$750

$3,340

Work in process inventory (WIP)

$522

$2,685

Finished goods inventory (FGI)

$469

$1,925

What are the days of supply for Dell?

Round your answer to 1 decimal place and do not list the units (for example, answer like this: 24.3)
Answer #1

Total inventory cost = (PARTS + WIP + FGI) = 750 + 522 + 469 = $ 1741

Cost of goods sold = $ 10,105

Days of supply = inventory cost * 365 / cost of goods sold.

Days of supply = (1741 * 365) / 10,105 = 62.9